# Collaborative Deterministic–Probabilistic Forecasting for Real-World Spatiotemporal Systems

### v4

@@ C.3 Experimental Configuration
In our experiment, for our model, we set the training maximum epoch for both the deterministic model and the diffusion model to 50, with early stopping based on patience of 5 for both models. For the diffusion model, we set the validation set sampling number to 3, and the average metric computed over these samples is used as the criterion for early stopping. For the baseline models, we set the maximum training epoch to 100 and the early stopping patience also to 5. We set the number of samples to 50 for computing the experimental results presented in the paper. For the denoising network architecture, we adopt a lightweight variant of the MLP-based STID [49]. Specifically, we set the number of encoder layers to 8 and the embedding dimension to 128. The diffusion model employs a maximum of 50 diffusion steps, using a linear noise schedule with $\beta_{1}=0.0001$ and $\beta_{N}=0.5$. During training, we set the initial learning rate to 0.001, and after 20 epochs, we adjust it to 4e-4. We use the Adam optimizer with a weight decay of 1e-6. All experiments are conducted with fixed random seeds. Models with lower GPU memory demands are run on NVIDIA TITAN Xp (12GB GDDR5X) and NVIDIA GeForce RTX 4090 (24GB GDDR6X) GPUs under a Linux environment. For the DYffusion [47] baseline, which requires substantially more resources, training is performed on NVIDIA A100 (80GB HBM2e) and A800 (40GB HBM2e).
C.4 Geographic Extent of the ENSO Region
@@
C.5 Additional Experimental Results
Table 6: Short-term forecasting results in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and ‘-’ denotes results that are not applicable.
@@ C.5 Additional Experimental Results
Table 7: Short-term forecasting results in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and ‘-’ denotes results that are not applicable.
@@
Table 8: Long-term forecasting results in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and ‘-’ denotes results that are not applicable.
@@
Table 9: Long-term forecasting results in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and ‘-’ denotes results that are not applicable.
@@
C.5.1 Analysis of Distribution Alignment.

### v5

@@ E.3 Experimental Configuration
In our experiment, for our model, we set the training maximum epoch for both the deterministic model and the diffusion model to 50, with early stopping based on a patience of 5 for both models. For the diffusion model, we set the validation set sampling number to 3, and the average metric computed over these samples is used as the criterion for early stopping. For the baseline models, we set the maximum training epoch to 100 and the early stopping patience also to 5. We set the number of samples to 50 for computing the experimental results presented in the paper. For the denoising network architecture, we adopt a lightweight variant of the MLP-based STID (Shao et al., 2022). Specifically, we set the number of encoder layers to 8 and the embedding dimension to 128. The diffusion model employs a maximum of 50 diffusion steps, using a linear noise schedule with $\beta_{1}=0.0001$ and $\beta_{N}=0.5$. During training, we set the initial learning rate to 0.001, and after 20 epochs, we adjust it to 4e-4. We use the Adam optimizer with a weight decay of 1e-6. All experiments are conducted with fixed random seeds. Models with lower GPU memory demands are run on NVIDIA TITAN Xp (12GB GDDR5X) and NVIDIA GeForce RTX 4090 (24GB GDDR6X) GPUs under a Linux environment. For the DYffusion (Rühling Cachay et al., 2023) baseline, which requires substantially more resources, training is performed on NVIDIA A100 (80GB HBM2e) and A800 (40GB HBM2e).
E.4 Geographic Extent of the ENSO Region
@@
E.5 Additional Experimental Results
Table 6: Short-term forecasting results in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and “-” denotes results that are not applicable.
@@ E.5 Additional Experimental Results
Table 7: Short-term forecasting results in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and “-” denotes results that are not applicable.
@@
Table 8: Long-term forecasting results in terms of CRPS, QICE, and IS. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and “-” denotes results that are not applicable.
@@
Table 9: Long-term forecasting results in terms of MAE and RMSE. Bold indicates the best performance, while underlining denotes the second-best. DYffusion is limited to grid-format data, and “-” denotes results that are not applicable.
@@
Table 10: Comparison of training and inference time on the MobileSH dataset.
Model Train Time Inference Time
D3VAE 3min 27s 2min 15s
DiffSTG 24min 16s 18min 38s
TimeGrad 5min 2min
CSDI 48min 40s 38min 49s
DyDiffusion 33h 3h
CoST 2min 50s
Table 11: Long-term forecasting performance comparison of TMDM on ETTh1 and ETTh2 Datasets.
Model EETh1 EETh2
CRPS QICE IS CRPS QICE IS
TMDM 0.395 0.041 4.8 0.196 0.018 2.2
CoST 0.311 0.007 1.6 0.109 0.007 0.78
E.5.1 Analysis of Distribution Alignment.